# Keeping Medical AI Healthy: A Review of Detection and Correction Methods for System Degradation

### v1

@@ 8.1.2 Test-Time Domain Adaptation (TTDA)
TTDA [95] refers to a set of methods that adapt models on-the-fly during inference using only unlabeled target data observed at test time. Unlike UDA, TTDA does not require access to target data during training and operates under more constrained but realistic conditions.
TTDA has been applied in cross-site medical image classification and segmentation tasks, demonstrating adaptability to distribution shifts, particularly covariate shifts [96, 97]. Recently, TTDA has also been extended to LLMs in biomedical NLP. [98] introduces TTDA to improve LLM performance on tasks such as biomedical question answering, showing that even large pretrained models can benefit from dynamic adjustment to target-domain inputs at inference time.
@@ 8.2.1 Re-Training
The study [27] examines the effects of event-driven shifts (e.g., during COVID-19) on sepsis prediction models. Simulation results show that retrained models, especially XGBoost, outperformed static models, highlighting the importance of regular retraining to maintain performance. [99] explores how feedback loops in healthcare can degrade models over time. Using MIMIC-IV data, they showed that repeated updates may increase false positive rates, even when AUC appears stable. Their findings indicate that retraining on the full historical data is more robust than incremental updates using only recent data.
Despite its effectiveness, retraining is resource-intensive, requiring quality labels and full retraining-validation cycles. These challenges may limit feasibility in fast-paced clinical environments.
@@ 8.2.2 Fine-Tuning (Transfer Learning)
It adapts a pretrained model to a new domain by updating all or part of its parameters using some target data. It is particularly effective when source and target domains share structural similarities but differ in characteristics (e.g., patient demographics /imaging protocols).
@@
Fine-tuning offers a cost-effective and flexible correction method, especially when labeled data are scarce but pretrained models offer transferable features. It is well-suited for cross-site clinical model deployment.
8.3 Continual Learning (CL)
CL [101, 102] enables AI models to adapt to new data over time while preserving prior knowledge. Unlike static models, CL supports dynamic updates, making it well-suited for healthcare settings where disease patterns, practices, and demographics evolve continuously.
[103] highlights the potential of CL in diagnostic imaging, prediction, and treatment planning, while also noting challenges such as catastrophic forgetting, computational overhead, and regulatory concerns. They stress the importance of robust validation and adaptive deployment strategies. [100] applies drift-triggered CL in a hospital setting, detecting temporal shifts and retraining on updated data. This approach successfully maintains model performance under changing clinical conditions. [104] proposes a lifelong learning framework (a type of CL) using knowledge distillation to address calibration drift. Tested on four cancer datasets, the method outperformed standard retraining by preserving both calibration and discrimination over time.
8.4 Calibration Correction
Calibration correction [105] adjusts a model’s predicted probabilities to better reflect true outcome rates in the target population. Unlike retraining methods, it modifies outputs directly, maintaining accurate risk estimates. This is important in healthcare, where decisions often depend on calibrated thresholds.
[106] categorizes calibration update strategies into: (1) simple coefficient updating (adjusting intercepts and slopes), (2) meta-model updating (combining existing models), and (3) dynamic updating (continual refinement with new data). Using cardiac surgery data, they show that adaptive updates improve reliability in clinical settings.
Recent studies explore label-free approaches. [88] uses cohort-level features to detect and correct calibration shifts without labels. Similarly, [107] aligns prediction distributions across imaging datasets to mitigate acquisition shifts, preserving performance in tasks like breast cancer detection without labeled target data.
@@ 8.5 Adaptive Update Strategies
Update strategies guide when and how to apply corrections in response to performance degradation. Unlike individual correction techniques, they provide a broader decision-making framework essential for managing real-world shifts that vary in scale and frequency. Effective strategies balance timely intervention with model stability. Key questions include:
  • When to update? Only if the shift significantly impacts performance.
  • How to update? Options include recalibration, partial fine-tuning, or full retraining.
  • How often to update? Avoid excessive or delayed updates to maintain reliability.
@@ 9.1 Datasets
Medical AI monitoring remains challenging due to the scarcity of publicly available datasets and standardized benchmarks, as most studies rely on proprietary hospital data. Advancing the field requires open datasets that capture both “natural shifts” (e.g., those arising real-world events like the COVID-19 pandemic) and “synthetic shifts” (e.g., engineered changes in demographics or image quality) to rigorously evaluate model robustness. The COVID-19 pandemic highlights the need to monitor data shift in medical AI. For instance, the University of Michigan Hospital deactivated a sepsis-alerting model in April 2020 after demographic shifts triggered spurious alerts [25]. Such shifts can disrupt clinical variable relationships, degrade model performance, and increase patient risks.
@@ 9.3.1 Stanford WILDS/Wild-Time Benchmarks
The Stanford WILDS benchmark [126] [^1] https://wilds.stanford.edu/ evaluates model robustness under real-world distribution shifts across domains like healthcare and ecology. Results show significant performance drops from in-distribution to out-of-distribution settings, emphasizing the challenge of robust deployment. Extending this, Wild-Time [127] [^2] https://wild-time.github.io focuses on temporal shifts using five real-world datasets, including MIMIC-IV for clinical prediction tasks. Wild-Time reports  20% performance drops over time, highlighting the difficulty of achieving temporal robustness.
@@ 9.3.2 TableShift Benchmark
TableShift [128] [^3] https://tableshift.org is a benchmark designed to evaluate distribution shift robustness of machine learning on tabular data, an area that, despite its widespread use, has lacked standardized benchmarks compared to text and vision. This is especially important in healthcare, where structured EHR data is common. TableShift includes clinical tasks such as ICU mortality, length of stay, and sepsis prediction.
It supports a range of tabular models, including state-of-the-art architectures like FT-Transformer [129] and Tabular ResNet [130], and provides Python APIs for streamlined data processing and evaluation.
@@ 9.3.3 Concept Drift Detection Benchmark
In A Large-Scale Comparison of Concept Drift Detectors [131], 14 drift detection methods are evaluated on synthetic datasets with abrupt and gradual shifts, using Naïve Bayes and Hoeffding Tree as the two basic classifiers. It evaluates both classification performance and drift detection accuracy, and analyze their correlation. One interesting finding is that precise detection of all drifts may not always lead to better outcomes.
@@ 10.1.1 Missing or Delayed Ground-Truth Labels for AI Monitoring
Ideally, real-time monitoring of AI models allows for the early detection of performance degradation, enabling healthcare staff to respond quickly. However, obtaining ground-truth labels, the gold standard for performance evaluation, is often delayed or infeasible in clinical settings. Label generation typically requires significant time, labor, and cost, such as clinician annotation of chest X-rays for pulmonary diagnoses. Moreover, definitive clinical outcomes, such as mortality or disease profession, often become available only after significant delays. These limitations impede timely performance assessment and delay corrective interventions, particularly for predictive and diagnostic models targeting long-term outcomes.
@@ 10.1.2 Human-Centric Requirement Monitoring
Health AI monitoring is a rapidly evolving field, traditionally focused on performance metrics like diagnostic accuracy, F1-score, and AUC. However, recent research underscores the need to address human-centric concerns such as privacy, fairness, and alignment with human values [132]. Performance metrics alone are insufficient. For example, a system may show stable accuracy overall while underperforming for certain ethnic groups. To ensure equitable and trustworthy AI, monitoring must also detect and address such disparities. Future efforts should incorporate tools like bias and privacy detectors to complement performance monitors and support both technical reliability and societal accountability.
@@ 10.1.3 Stability-Plasticity Dilemma
A key challenge in updating AI models after performance degradation is the stability-plasticity dilemma [133], the trade-off between preserving prior knowledge (stability) and adapting to new information (plasticity). Stability: Models must retain learned decision boundaries to maintain consistent clinical support. Frequent updates risk catastrophic forgetting [30, 134], compromising previously reliable predictions. Plasticity: Models must adapt to new medical knowledge, shifts in disease prevalence, and evolving clinical guidelines. Without adaptability, they become outdated and misaligned with current patient needs. Balancing these forces remains difficult in practice. Excessive retraining can destabilize models, while infrequent updates lead to performance decay. Robust adaptive strategies are essential to keep AI models both reliable and relevant.
@@ 10.1.4 System Complexity and Maintenance Challenges
Real-world AI systems are increasingly complex, comprising feedback loops, dynamic data pipelines, and interdependent modules. Unlike rule-based software, their behavior is data-driven, making them sensitive to input shifts and operational changes. Such complexity increases the risk of silent degradation. Feedback loops may reinforce bias, and changes in one component can unintentionally destabilize others, accumulating what is often referred to as technical debt [135]. These challenges are magnified in large models like LLMs and VLMs, whose black-box architectures and emergent behaviors (unintended capabilities arising post-training) make them difficult to interpret or monitor [136].
Manual oversight is no longer sufficient. Ensuring long-term safety and reliability demands automated, adaptive monitoring systems that can detect hidden failures, track behavioral shifts, and improve model transparency.
@@ 10.2.1 Label-Free Estimation
In medical AI, ground-truth labels are often delayed, costly, or impractical to obtain, making traditional retrospective validation insufficient for real-time model assessment. Label-free performance estimation offers a promising alternative, allowing models to self-assess without labeled target data.
@@ 10.2.5 Monitoring Large Language and Vision-Language Models
LLMs and VLMs are increasingly used in healthcare, but their massive scale and black-box nature raise concerns about safety and reliability in clinical settings. Recent studies reveal risks of behavioral drift. For example, [21] find that GPT-4’s instruction-following ability declined over time, highlighting the need for ongoing monitoring. Performance degradation in LLMs/VLMs may result from knowledge staleness, hallucination, catastrophic forgetting, or few-shot performance decay. Empirical evidence [19, 20, 40] shows that GPT-4’s diagnostic accuracy dropped within months across both text and image tasks. As these models play a growing role in clinical decision-making, monitoring their performance is a critical and promising research direction.
@@ 11 Conclusion
This survey presents a comprehensive framework for understanding and addressing AI performance degradation in healthcare, covering data and model drift detection, root cause analysis, and correction strategies. We also review representative open-source datasets, tools, and benchmarks that support medical AI monitoring. We highlight key challenges, such as delayed ground truth, fairness, the stability-plasticity trade-off, and system complexity, and emphasize future directions including label-free performance estimation, shift impact assessment, root cause analysis, synthetic data use, and continuous monitoring of large-scale models to ensure safety and reliability.
Acknowledgments

### v2

@@ 8.1.2 Test-Time Domain Adaptation (TTDA)
TTDA [101] refers to the methods that adapt models on-the-fly during inference using only unlabeled target data observed at test time. Unlike UDA, TTDA does not require access to target data during training and operates under more constrained but realistic conditions.
TTDA has been applied in cross-site medical image classification and segmentation tasks, demonstrating adaptability to distribution shifts, particularly covariate shifts [102, 103, 104, 105]. Recently, TTDA has also been extended to LLMs in biomedical NLP. [106] introduces TTDA to improve LLM performance on tasks such as biomedical question answering, showing that even large pretrained models can benefit from dynamic adjustment to target-domain inputs at inference time.
@@ 8.2.1 Re-Training
The study [33] examines the effects of event-driven shifts (e.g., during COVID-19) on sepsis prediction models. Simulation results show that retrained models, especially XGBoost, outperformed static models, highlighting the importance of regular retraining to maintain performance. [107] explores how feedback loops in healthcare can degrade models over time. Using MIMIC-IV data, they show that repeated updates may increase false positive rates, even when AUC appears stable. Their findings indicate that retraining on the full historical data is more robust than incremental updates using only recent data.
Despite its effectiveness, retraining is resource-intensive, requiring quality labels and full validation cycles, which may limit its feasibility in fast-paced clinical settings.
@@ 8.2.2 Fine-Tuning (Transfer Learning)
It adapts a pretrained AI model to a new domain by updating all or part of its parameters using some target data. It is particularly effective when source and target domains share structural similarities but differ in characteristics (e.g., patient demographics or imaging protocols).
@@
Fine-tuning offers a cost-effective and flexible correction method, especially when labeled data are scarce but pretrained models offer transferable features.
@@
8.3 Continual Learning (CL)
CL [122, 123] enables models to adapt to new data while retaining prior knowledge, offering dynamic updates suited to evolving healthcare environments (e.g., disease patterns, practices, and demographics).
[124] highlights the potential of CL in diagnostic imaging, prediction, and treatment planning, while also noting challenges such as catastrophic forgetting, computational overhead, and regulatory concerns. [108] applies drift-triggered CL in a hospital setting, detecting temporal shifts and retraining on updated data. This approach successfully maintains model performance under changing clinical conditions. [125] proposes a lifelong learning framework using knowledge distillation to mitigate calibration drift, outperforming standard retraining on four cancer datasets by maintaining both calibration and discrimination over time.
8.4 Calibration Correction
Calibration correction [126] adjusts a model’s predicted probabilities to better reflect true outcome rates. Unlike retraining methods, it modifies outputs directly, maintaining accurate risk estimates. This is important in healthcare, where decisions often depend on calibrated thresholds. Su et al.[127] categorize calibration updates into 1) coefficient adjustment, 2) meta-model combination, and 3) dynamic updating with new data. Using cardiac surgery data, they show that adaptive updates improve reliability in clinical settings.
Recent studies explore label-free approaches. [92] uses cohort-level features to detect and correct calibration shifts without labels. Similarly, [116] aligns prediction distributions across imaging datasets to mitigate acquisition shifts, preserving performance in tasks like breast cancer detection.
@@ 8.5 Adaptive Update Strategies
Update strategies define when and how to apply corrections for performance degradation, offering a broader framework for managing shifts of varying scale and frequency. Effective strategies balance timely intervention with model stability. Key questions include: 1) When to update? Only if the shift significantly impacts performance; 2) How to update? Options include recalibration, partial fine-tuning, or full retraining. 3) How often to update? Avoid excessive or delayed updates to maintain reliability.
@@
To summarize the above methods with the qualitative discussion, Table 1 presents representative approaches within the Detection-Diagnosis-Correction framework, highlighting their main prerequisites and relative computational cost.
@@ 9.1 Datasets
Medical AI monitoring remains challenging due to the scarcity of publicly available datasets and standardized benchmarks, as most studies rely on proprietary hospital data. Advancing the field requires open datasets that capture both “natural shifts” (e.g., those arising from real-world events like the COVID-19 pandemic) and “synthetic shifts” (e.g., engineered changes in demographics or image quality) to evaluate model robustness. The COVID-19 pandemic highlights the need to monitor medical AI. For instance, the University of Michigan Hospital deactivated a sepsis-alerting model in April 2020 after demographic shifts triggered spurious alerts [31]. Such shifts can disrupt clinical variable relationships, degrade model performance, and increase patient risks.
@@ 9.3.1 Stanford WILDS/Wild-Time Benchmarks
The Stanford WILDS benchmark [138] [^1] https://wilds.stanford.edu/ evaluates model robustness under real-world distribution shifts across domains like healthcare and ecology. Results show significant performance drops from in- to out-of-distribution settings, underscoring the challenge of robust deployment. Extending this, Wild-Time [139] [^2] https://wild-time.github.io focuses on temporal drift using five real-world datasets, including MIMIC-IV for clinical prediction tasks. Wild-Time reports around 20% performance drops of AI models over time, highlighting the difficulty of achieving temporal robustness.
@@ 9.3.2 TableShift Benchmark
TableShift [140] [^3] https://tableshift.org benchmarks distribution-shift robustness in tabular data, an area long lacking standardized evaluation compared to text and vision. This is critical in healthcare, where structured EHR data are common. It includes tasks such as ICU mortality, length of stay, and sepsis prediction, supports models like FT-Transformer [141] and Tabular ResNet [142].
@@ 9.3.3 Concept Drift Detection Benchmark
In the Large-Scale Comparison of Concept Drift Detectors [143], 14 drift detection methods are evaluated on synthetic datasets with abrupt and gradual shifts, using Naïve Bayes and Hoeffding Tree as the two basic classifiers. It evaluates both classification performance and drift detection accuracy, and analyze their correlation. One interesting finding is that precise detection of all drifts may not always lead to better outcomes.
9.3.4 Discussion: Benchmarking and the Evaluation Gap
While several public datasets support data shift and OOD research (e.g., WILDS, MedShift, OpenOOD), most benchmarks remain general-purpose and do not fully capture the clinical complexity of healthcare settings. As recently emphasized by [144], biomedical machine learning faces a “benchmarking crisis”, where performance is often reported without rigorous or standardized evaluation across institutions, populations, and acquisition settings. This limitation extends to AI monitoring, where few benchmark suites simulate longitudinal or multi-site deployment. Establishing standardized, clinically grounded benchmarks for AI monitoring is therefore an urgent priority for enabling fair, reproducible, and thorough evaluation.
@@ 10.1.1 Missing or Delayed Ground-Truth Labels for AI Monitoring
Real-time monitoring enables early detection of AI performance degradation, allowing timely clinical response. However, ground-truth labels, the gold standard for evaluation, are often delayed or infeasible in clinical settings. Label generation typically requires expert annotation (e.g., chest X-ray annotations for pulmonary diagnoses) or long-term outcomes (e.g., mortality, disease progression). Such delays hinder timely assessment and correction, especially for predictive and diagnostic models.
@@ 10.1.2 Human-Centric Requirement Monitoring
Medical AI monitoring has traditionally emphasized performance metrics such as diagnostic accuracy. However, recent research underscores the need to address human-centric concerns such as privacy, fairness, and alignment with human values [145]. Performance metrics alone are insufficient. For example, a system may show stable accuracy overall while underperforming for certain ethnic groups. To ensure equitable and trustworthy AI, monitoring must also detect and address such disparities. Future efforts should incorporate tools like bias and privacy detectors to complement performance monitors and support both technical reliability and societal accountability.
@@ 10.1.3 Stability-Plasticity Dilemma
A key challenge in updating AI models after degradation is the stability-plasticity dilemma [146], the trade-off between retaining prior knowledge (stability) and adapting to new information (plasticity). Stability ensures consistent clinical support but insufficient updates cause obsolescence. Plasticity enables adaptation to evolving medical knowledge, disease patterns, and clinical guidelines, but frequent updates risk catastrophic forgetting [36, 147]. Balancing these forces is difficult: excessive update destabilizes models, while infrequent updates cause performance decay. Effective adaptive strategies are crucial to keep AI systems both reliable and relevant.
@@ 10.1.4 System Complexity and Maintenance Challenges
Real-world AI systems are increasingly complex, featuring feedback loops, dynamic data pipelines, and interdependent modules. Unlike rule-based software, their data-driven nature makes them sensitive to input shifts and operational changes. This complexity heightens the risk of silent degradation: feedback loops can amplify bias, and changes in one component may destabilize others, accumulating technical debt [148]. These challenges are further magnified in large models such as LLMs and VLMs, whose black-box architectures and emergent behaviors make interpretation and monitoring difficult [149].
10.1.5 Prompt-Injection and Retrieval Drift in LLM-Based Systems
For LLM-based medical systems, new forms of drift emerge beyond traditional data or model shifts and require specialized monitoring strategies. Prompt injection can be viewed as an input-level anomaly in which malicious or unintended instructions alter model behavior. Such anomalies can be detected by monitoring deviations in prompt structure, intent, or embedding similarity against validated templates. Retrieval drift refers to the gradual evolution of external knowledge sources in retrieval-augmented generation (RAG) pipelines, leading to inconsistent or outdated information. It can be monitored through periodic validation of retrieved content. Both forms of drift introduce new challenges that call for further research.
10.1.6 Multimodal AI Monitoring and Missing Modality Challenges
As medical AI systems increasingly integrate diverse data types (e.g., imaging, EHRs, genomics, and text), monitoring multimodal models introduces additional complexity. Unlike unimodal systems, multimodal AI must maintain consistency across heterogeneous input spaces, making it sensitive to cross-modal drift (e.g., distribution shifts in one modality that alter joint representations) and missing modality scenarios that frequently occur in real-world practice. Recent work, such as MultiOOD [80], highlights scalable OOD detection across multiple modalities. Addressing such challenges requires monitoring frameworks that jointly assess modality-specific and cross-modal reliability, with correction strategies that adapt to missing or degraded modalities at inference.
@@ 10.2.1 Label-Free Estimation
In medical AI, ground-truth labels are often delayed or costly, limiting real-time monitoring. Label-free performance estimation offers an alternative, enabling models to self-assess without labeled target data.
@@ 10.2.5 Monitoring Large Language and Vision-Language Models
LLMs and VLMs are increasingly used in healthcare, but their massive scale and black-box nature raise concerns about safety and reliability in clinical settings. Recent studies reveal risks of behavioral drift. For example, [25] finds that GPT-4’s instruction-following ability declined over time, highlighting the need for ongoing monitoring. Performance degradation in LLMs/VLMs may result from knowledge staleness, hallucination, catastrophic forgetting, or few-shot performance decay. Empirical evidence [23, 24, 46] shows that GPT-4’s diagnostic accuracy dropped within months across both text and image tasks. As these models play a growing role in clinical decision-making, monitoring their performance is a critical and promising research direction.
@@ 11 Conclusion
This survey presents a comprehensive framework for understanding and addressing AI performance degradation in healthcare, covering data and model drift detection, root cause analysis, and correction strategies. We also review representative open-source datasets, tools, and benchmarks that support medical AI monitoring. We highlight key challenges, such as delayed labels, fairness, the stability-plasticity trade-off, system complexity and multimodality, and emphasize future directions including label-free performance estimation, root cause analysis, synthetic data use, and continuous monitoring of large-scale models to ensure safety and reliability.
Acknowledgements